Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Sun Feb 14 08:01:33 MET 1999) Minor failures
Lost few scans at the beginning

Westerbork (Tony Foley at Sun Feb 14 16:28:59 MET 1999) Failure
Only recorded to from 00:00 to 01:02UT as we had a tape break at 13000ft on the third pass. Problems rewinding and removing tape. The reason for the tape breaking is not clear, as it had gone through prepass OK a few days ago. We used the tied array (8MHz of the 16MHz).

Jodrell (Alastair Gunn at Sun Feb 14 18:16:36 MET 1999) Success
Communication with IF distributor at start of experiment but problem was resolved before start of data acquisition. No other known problems.

Onsala (A. Polatidis at Mon Feb 15 09:47:27 MET 1999) Success

Torun (K.Borkowski at Mon Feb 15 13:16:46 MET 1999) Success
No known problems.

Noto (C.Trigilio at Tue Feb 16 18:24:38 MET 1999) Minor failures
Lost first 2 scans due to a software problem during field-system re-configuration (from MK3 to VLBA mode). Fist valid data at 00:45 (source 0859+470).

Effelsberg (dgraham at Mon Feb 22 15:19:57 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:27:20 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: v030k.antab. Read the comments in the file to get further information. (Missed Sh log file)